OpenStudy (anonymous):

@Maria195 plz help me

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

Ok so do you know the linaer equation?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

plug in the points for x and y

OpenStudy (anonymous):

and the slope of -1.2 and solve for b.

OpenStudy (anonymous):

It's from 3 to 5 because the horizontal line goes from 3 to 5 on the x-axis.

OpenStudy (anonymous):

So you get 8=-1.2(7)+b

OpenStudy (anonymous):

wait so its y=blank x+blank

OpenStudy (anonymous):

simply this

OpenStudy (anonymous):

mutiply -1.2 by 7 and you get -8.4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

by doing this you get 8=8.4+b so b must equal -.4

OpenStudy (anonymous):

blank 1 =-1.2 and blank 2=-.4